# Learning Program Behavioral Models from Synthesized Input-Output Pairs

### v1

@@ Abstract.
We introduce Modelizer—a novel framework that, given a black-box program, learns a model from its input/output behavior using neural machine translation. The resulting model mocks the original program: Given an input, the model predicts the output that would have been produced by the program. However, the model is also reversible—that is, the model can predict the input that would have produced a given output. Finally, the model is differentiable and can be efficiently restricted to predict only a certain aspect of the program behavior. Modelizer uses grammars to synthesize inputs and to parse the resulting outputs, allowing it to learn sequence-to-sequence associations between token streams. Other than input and output grammars, Modelizer only requires the ability to execute the program. The resulting models are small, requiring fewer than 6.3 million parameters for languages such as Markdown or HTML; and they are accurate, achieving up to 95.4% accuracy and a BLEU score of 0.98 with standard error 0.04 in mocking real-world applications. We foresee several applications of these models, especially as the output of the program can be any aspect of program behavior. Besides mocking and predicting program behavior, the model can also synthesize inputs that are likely to produce a particular behavior, such as failures or coverage.
@@ 1. Introduction
Modelizer learns the model from the PUT by executing it again and again, observing how its inputs translate to its outputs. However, like all machine learning models, Modelizer depends on good training. This raises two central problems, for each of which Modelizer has a solution:
Synthesizing inputs.:
The first central problem thus is the availability of input data for training. In Software Engineering, the field of test generation discusses how to obtain input data for a program. The test data must be valid, such that it gets accepted by the PUT, as well as diverse, such that a wide variety of input features is covered. In a black-box setting, obtaining valid and diverse input data requires a dedicated generator or producer, leveraging knowledge about the input language of the program.
Tokenization.:
The second problem is obtaining features from inputs and outputs. Modelizer uses a sequence-to-sequence translation model to translate token sequences from inputs into outputs. Hence, we need a decomposition of inputs and outputs into individual tokens, so we can learn how token sequences map into each other. While this is a mostly solved problem for natural languages, tokenizing formal inputs correctly is a challenge, notably as we want the tokenizer to be general-purpose, and suitable for all sorts of input and output languages.
Modelizer addresses both problems by leveraging grammars for both input and output language. The input grammar is used to produce syntactically valid and diverse inputs; it can also parse existing inputs into derivation trees. In both cases, we know the structure of the input. The output grammar is used to parse the output, again delivering a tree structure allowing decomposition into tokens. Using the input grammar as a producer, we can synthesize valid and diverse input data automatically and thus obtain as many input-output pairs from the PUT as we want. In cases when output language grammar is unavailable, the program output can be decomposed by unsupervised tokenizers. Modelizer distinguishes structure from content in the decomposed sequences and can abstract content by replacing them with placeholders to improve the learning quality.
@@
These scenarios only scratch the surface of possible usages. In this paper, however, we focus on the mock and reverse predictor scenarios—that is, predicting outputs from inputs and vice versa. We evaluate Modelizer across a range of challenging PUTs and achieve excellent accuracy. Notably, the models produced by Modelizer showed high accuracy on non-trivial programs despite having a small number of neurons (and thus efficient to learn). On the Pandoc markup converter, for instance, Modelizer achieved over 90% accuracy in converting Markdown to HTML and vice-versa with a BLEU score of 0.71. While complex computations as performed by or many commercial programs remain out of reach for our current methods, our experiments indicate that increased model and training sizes may make such learning possible in the future.
@@
To the best of our knowledge, Modelizer is the first approach to automatically learn sequence-to-sequence models from arbitrary programs.
@@
Modelizer uses grammars to produce inputs (and parse given inputs), and can thus learn from an unbounded number of training inputs and resulting outputs.
Abstract placeholders for tokenizers.:
Tokenizers that implement the suggested interface can automatically replace the real content with abstract placeholders, thus focusing on translating text structure rather than text content.
@@
The remainder of this paper is organized as follows: Section 2 describes in detail how Modelizer works, including input generation, dataset preprocessing, model learning, and model deployment. Section 3 covers the implementation aspects of Modelizer. In Section 4, we evaluate Modelizer on a variety of text-to-text translators, such as Markdown to HTML, query languages, and more. Section 5 discusses the limitations of Modelizer. After discussing related work (Section 6), Section 7 closes with a conclusion and future work. Modelizer and all experimental data are available as open source; see Section 7 for details.
@@ 2. Approach
Figure 2. How Modelizer works. Modelizer tests program with synthesized inputs and automatically learns reversible program behavior model from extracted input-output pairs.
@@
How Modelizer works. Modelizer tests program with synthesized inputs and automatically learns reversible program behavior model from extracted input-output pairs.
Modelizer is a generic framework for learning program behavior. It uses synthesized data to automatically learn reversible generative models that can replicate the behavior of real-world programs with high accuracy. Figure 2 shows a general overview of the framework. Given the input specification for the selected PUT, Modelizer automatically generates the required number of inputs and passes them to the PUT for processing. Once the PUT produces the corresponding output, Modelizer extracts these input-output pairs and uses them to train a reversible neural machine translation model. The trained model can be further deployed on modern computer architectures. We have automated program behavior mocking using the following steps: Input Generation, Dataset Preprocessing, Model Learning, and Model Deployment. The following sections will cover each step in detail.
@@ 2.1. Input Generation
While developing our framework, we aimed to make it interoperable with different types of programs. Even though all programs are written in different programming languages, differ in implementation logic, and target different hardware architectures and operating systems, most programs are made to process input coming from users or other programs. Thus, we have decided to learn program behavior by training our behavior mocking models from the program inputs and processed outputs. Of course, such a strategy will enforce certain limitations:
  • The program behavior should be deterministic. At every execution, the program should produce the same output for the given input.
  • We will not be able to model the behavior of programs that process extremely long inputs or outputs at once. The computational complexity, memory, and time constraints scale quadratically with the input-output length.
  • Learning accurate inverse models that predict inputs for the given output is challenging when PUT computes multiple inputs to the same output. An intermediate representation for inputs could assist in such cases, which leads to context loss.
To learn an accurate model, we need sufficiently many diverse input-output pairs. Since public test datasets are typically biased towards commonly used features, and we do not know in advance how many inputs are required for our approach to learn the behavior of an unseen program, Modelizer assumes that inputs are automatically generated.
@@ 2.1.1. Input Generation with Grammars
To cover the entire input space (and hence increase our chances to also cover the behavior and output spaces), we use grammar-based input producers (Zeller et al., 2023). These encode input specifications as context-free grammars (CFG) (Chomsky, 1956)—a finite set of recursive string generation rules. For some input formats, such CFGs already exist and are publicly available. For example, the Grammar Zoo (Zaytsev, 2015) repository contains definitions of 1756 grammars of mainstream programming languages mined from standard documentation or community-created resources.
Figure 3. Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification
@@
Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification
The first major component of the Modelizer framework is the Input Generator Module. This module functions as a data processing pipeline, which is sketched in Figure 3. It contains four steps: (1) Synthesis, (2) Post-Processing, (3) Hashing, and (4) Validation.
At the synthesis step, the input producer synthesizes new input samples according to the rules encoded in grammar. Then, the synthesized samples get refined to better align with real-world data. Such a refinement helps to map synthesized input features to real-world data features without increasing the size of grammars and, thus, simplifies the model learning process. The refined samples are then checked for uniqueness by computing a hash of the synthesized sample. If the hash value is not present in the Hashes Dataset, the sample is passed to the PUT for further processing. Otherwise, all steps are repeated till a unique sample is generated. The PUT produces the corresponding output, which is saved in a persistent storage along with the input. The collected input-output pairs are later used to train a program behavior model.
@@
14<code_block> ::= ‘<text>‘ | ‘‘‘<text>‘‘‘ | "<empty_space><text>"
15<quoted_text> ::= ’<text>’ | "<text>"
16<text> ::= TEXT
17<url> ::= URL
@@
Let us reuse our handwritten Markdown grammar (LABEL:lst:markdown-grammar) to demonstrate the synthetic input generation process. We will follow the steps that Modelizer takes to produce inputs. You can find the high-level implementation of the input generation pipeline written in Python in LABEL:lst:generate. Once the input producer is configured, it is ready to synthesize as many inputs as needed (Step 1). Below we give an example of synthesized input that was generated by the producer:
@@
1def generate_data(grammar, min_terminals, max_terminals, total_elements, program_handler):
@@
5 fuzzer = GrammarFuzzer(grammar, min_terminals, max_terminals)
@@
The producer has synthesized one top-level Markdown element, which is a paragraph block. The paragraph block was expanded to a single paragraph text element which respectively contains a sequence of formatted text, hyperlink, and codeblock elements. One of the formatted text elements was further expanded to a bold text element. The synthesized input contains TEXT and URL placeholders instead of real character sequences. We have decided to define placeholder tokens for replacing the real-world data instead of encoding the whole language vocabulary into our grammar and synthesizing natural language sentences. We will elaborate more about placeholders in the follow-up section.
@@ 2.1.2. Placeholders
Synthesizing the data that is equivalent to real-world counterparts is a challenging task. Different programs can process the same data in various ways. This means that the grammar must be specified or refined according to program expectations. Depending on the complexity of input specification, the number of required input samples for efficient behavior learning can also become extremely large. We have figured out that for certain data types and programs, some fragments that are not affected or used by underlying implementation logic can be abstracted. For example, many text format converters only process the document structure, leaving the document content untouched. To benefit from such effects, we suggested and implemented two strategies in our input generation pipeline:
@@
Both strategies will help to capture large portions of possible input space with few samples and correctly map real-world inputs to generic representations and vice-versa during the inference phase. Instead of learning the behavior from real-world examples, we learn how the PUT processes placeholders.
@@
We assume that a TEXT placeholder can represent any character sequence in a real Markdown document. Similarly, a URL placeholder can substitute an arbitrary definition. By introducing such an abstraction layer, we can cover the whole space of structures of real-world documents with a relatively small number of synthesized samples.
While the input producer’s output is already well-formed and will pass the input parsing guards, it can be challenging to align real-world Markdown documents with the synthesized structure because it is impossible to distinguish the generated placeholders from each other. Therefore, we have included a post-processing phase in the input generation pipeline, for which we have implemented a generic Placeholder Processor. It must be initialized with the list of Placeholders that are encoded into selected grammar, like [TEXT, URL], and can automatically assign a unique identifier to every placeholder that it finds in the given sequence. Of course, it is possible to inject identifiers during the input generation. However, grammar-based input generators do not synthesize integers incrementally. They cannot guarantee that the generated identifier will be used in a synthesized sequence of characters only once. It is much faster and easier to generate a simple candidate sequence and then refine it using custom string-replacement or constraint-solving rules. The presence of unique identifiers improves the accuracy of both input-to-output and output-to-input translations, specifically in the cases when elements must be reordered during translation. During the tokenization process, which we will cover in the following sections, the respective data portions are assigned with a matching placeholder type that already includes a uniquely generated identifier. Optionally, the assignment of unique identifiers can be disabled.
Since grammar-based producers do not maintain a list of generated samples, we have implemented a sample validation procedure which first computes an HMAC-SHA-384 (Hansen and 3rd, 2011) value (Step 3.1) for the synthesized sample and then checks whether the computed hash value is stored in the Hashes Dataset (Step 3.2). If a computed hash value is absent in the dataset, the dataset gets updated with a new record (Step 4) and the synthesized input is passed to the PUT to compute output. Otherwise, the synthesized sample is dropped, and all input generation steps are repeated till the unique sample is synthesized. Corresponding input-output pairs are saved in a persistent storage for later processing. The collected input-output pairs are saved in a persistent storage for later processing.
@@ 2.2. Dataset Pre-processing
Machine translation is the task of automatically translating a text from one language to another. Neural machine translation models are the latest iteration of machine translation approaches that rely on deep neural networks to learn dependencies and patterns between sentences in the source and target languages. The ability to learn these dependencies from matching input-output pairs inspired us to reason about the program behavior modeling as a sequence-to-sequence translation task. In principle, the current generation of neural machine translation architectures that we use in our implementation can operate with the whole input at once. The computational complexity of these systems scales quadratically concerning the length of the input, and practical implementations do not follow this principle. Instead, text-to-text translations are performed by partitioning the whole input into sequences of small elements called tokens. Then, neural text translation models learn how to generate sequences of tokens in a target language by mapping the cross-token dependencies from a source language to a target language. Similarly, the synthesized input and computed output must be converted into batched tensors for the Learner to be able to process them. While text partitioning is a well-studied problem in the natural language processing domain (Manning et al., 2008), and there are already tools that can perform this task (Honnibal and Montani, 2017), we find the tokenization of program input and output to be a challenging task.
2.2.1. Pre-trained and Unsupervised Tokenizers
Instead of implementing tokenizers from scratch, one can potentially rely on existing generic tokenizers. For example, the HuggingFace repository (Wolf et al., 2020) provides access to many community-developed natural language tokenizers through the HuggingFace API. Such tokenizers cannot properly split the program data for several reasons. First, they ignore the data-specific or computation-specific features. Both factors impact the tokenization granularity, which is covered in the following paragraphs. Also, some tokenizers, like the pre-trained Llama3 tokenizer, modify the input by prepending or appending additional characters to the produced tokens. This can break the syntactical correctness when it comes to program data and can impact the quality of learned models.
While any input and output can be encoded as a sequence of characters, semantically it does not necessarily need to be textual data. Programs can accept or produce structural data. Such a structure contains knowledge that is relevant for behavior learning, but many existing tokenizers cannot distinguish the data structure from the content. Since, for certain computation tasks, like converting Markdown-to-HTML or SQL-to-KQL, the program is going to process the structure of the input while leaving content unchanged, a tokenizer needs to be able to map the fragments of the input to an abstract representation while processing. It is possible to augment the tokenization process of pre-trained tokenizers with additional pre and post-processing steps, where inputs can be transformed into semantically similar representations with placeholders replacing the fragments containing content. To build such a pipeline, the parser for the given data type has to exist and align with the input parsing rules of PUT.
Still, the program behavior learning approach is tokenizer-agnostic, and an unsupervised tokenizer like Google SentencePiece (Kudo and Richardson, 2018) is sufficient to quickly start the behavior learning process. However, the implementation of a data-specific tokenizer will help to further improve the quality of the learned model.

### v2

@@ Abstract.
We introduce Modelizer—a novel framework that, given a black-box program, learns a model from its input/output behavior using neural machine translation algorithms. The resulting model mocks the original program: Given an input, the model predicts the output that would have been produced by the program. However, the model is also reversible—that is, the model can predict the input that would have produced a given output. Finally, the model is differentiable and can be efficiently restricted to predict only a certain aspect of the program behavior. Modelizer uses grammars to synthesize and inputs and unsupervised tokenizers to decompose the resulting outputs, allowing it to learn sequence-to-sequence associations between token streams. Other than input grammars, Modelizer only requires the ability to execute the program. The resulting models are small, requiring fewer than 6.3 million parameters for languages such as Markdown or HTML; and they are accurate, achieving up to 95.4% accuracy and a BLEU score of 0.98 with standard error 0.04 in mocking real-world applications. As it learns from and predicts executions rather than code, Modelizer departs from the LLM-centric research trend, opening new opportunities for program-specific models that are fully tuned towards individual programs. Indeed, we foresee several applications of these models, especially as the output of the program can be any aspect of program behavior. Beyond mocking and predicting program behavior, the models can also synthesize inputs that are likely to produce a particular behavior, such as failures or coverage, thus assisting in program understanding and maintenance.
@@ 1. Introduction
Modelizer learns the model from the PUT by executing it again and again, observing how its inputs translate to its outputs. However, like all machine learning models, Modelizer depends on good training. This raises data availability problem, for which we need to provide a solution. In Software Engineering, the field of test generation discusses how to obtain input data for a program. The test data must be valid, such that it gets accepted by the PUT, as well as diverse, such that a wide variety of input features is covered. In a black-box setting, obtaining valid and diverse input data requires a dedicated generator or producer, leveraging knowledge about the input language of the program.
Modelizer addresses the data availability problem by synthesizing inputs with the help of grammars - formal languages for generating strings. The input grammar is used to produce syntactically valid and diverse inputs; it can also parse existing inputs into derivation trees. In both cases, we know the structure of the input. The output grammar can also be used to parse the output, again delivering a tree structure allowing decomposition into tokens; if the output language specification is unknown, the program output can be decomposed by automatically training a tokenizer from scratch using the SentencePiece (Kudo and Richardson, 2018) algorithm. Using the input grammar as a producer, we can synthesize valid and diverse input data automatically and thus obtain as many input-output pairs from the PUT as we want. Additionally, Modelizer distinguishes structure from content in the decomposed sequences and can abstract content by replacing them with placeholders to further improve the learning quality and efficiency.
@@
These scenarios only scratch the surface of possible usages. In this paper, however, we focus on the mock and reverse predictor scenarios—that is, predicting outputs from inputs and vice versa. We evaluate Modelizer across a range of challenging PUTs and achieve excellent accuracy. Notably, the models produced by Modelizer showed high accuracy on non-trivial programs despite having a small number of neurons (and thus efficient to learn) when compared to large language models (LLMs). For instance, on the Pandoc markup converter, Modelizer achieved over 90% accuracy in converting Markdown to HTML and vice-versa with a BLEU score of 0.71.
Despite the potential of Modelizer, we also identify several limitations:
  • Some complex programs remain out of reach for our current methods due to the computational complexity of the underlying learning algorithm which scales quadratically relative to the processed data length. The legnths of tokenized sequences with input-output data can vary significantly depending on the selected data decomposition and abstraction strategies. For example, intermediate abstraction layers can significantly shorten sequence lengths but can cause context loss, which will affect the prediction quality. Existing sequence-to-sequence translation models tend to make mistakes when predicting long sequences. Thus, lack of prior knowledge about the subject properties and hardware limits makes it hard to pre-specify all tasks where Modelizer will succeed.
  • We still require minor developer effort in configuring the Modelizer framework. Our experiments suggest that improving synthetic input generation is crucial for automating behavior extraction. Feedback from real program runs can already help to automate grammar tweaks, reconfigure fuzzers, and adjust learners’ hyperparameters. Continuous growth of the computational power will soon enable complete behavior extraction of any subjects.
@@
To the best of our knowledge, Modelizer is the first approach to automatically extract execution behavior from arbitrary programs by converting them into sequence-to-sequence translation models.
@@
Modelizer uses grammars to produce inputs (and parse given inputs), and, thus, can learn from an unbounded number of training inputs and resulting outputs.
@@
The remainder of this paper is organized as follows: Section 2 describes in detail how Modelizer works, including input generation, dataset preprocessing, model learning, and model deployment. Section 3 covers the implementation aspects of Modelizer. In Section 4, we evaluate Modelizer on a variety of text-to-text translators, such as Markdown to HTML, query languages, and more. Section 5 discusses the limitations of Modelizer; Section 6 details the related work. Modelizer and all experimental data are available as open source; see Section 7 for details.
@@ 2. Approach
Modelizer is a generic framework for learning program behavior. It uses synthesized data to automatically learn reversible generative models that can replicate the behavior of real-world programs with high accuracy. Figure 2 shows a general overview of the framework. Given the input specification of the selected PUT encoded as grammar, Modelizer automatically generates the required number of inputs and passes them to the PUT for processing. Once the PUT produces the corresponding output, Modelizer extracts these input-output pairs and uses them to train a reversible neural machine translation model. The trained model can be further deployed on modern computer architectures. We have automated program behavior mocking using the following steps: Input Generation, Dataset Preprocessing, Model Learning, and Model Deployment. The following sections will cover each step in detail.
Figure 2. How Modelizer works. Modelizer tests the program with synthesized inputs and automatically learns a reversible program behavior model from extracted input-output pairs.
@@
How Modelizer works. Modelizer tests the program with synthesized inputs and automatically learns a reversible program behavior model from extracted input-output pairs.
@@ 2.1. Input Generation
While developing the Modelizer framework, we aimed to make it interoperable with different types of programs. Even though all programs are written in different programming languages, differ in implementation logic, and target different hardware architectures and operating systems, most programs are made to process input coming from users or other programs. Thus, we have decided to learn the program behavior by training our behavior-mocking models from the program inputs and processed outputs.
To learn an accurate model, we need sufficiently many diverse input-output pairs. Since public test datasets are typically biased towards commonly used features, and we do not know in advance how many inputs are required for learning the behavior of an unseen program, Modelizer assumes that inputs are generated automatically.
@@ 2.1.1. Input Generation with Grammars
Figure 3. Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification. Input generation steps: (1) Synthesis, (2) Post-Processing, (3) Hashing, and (4) Validation.
@@
Input generation pipeline. Modelizer automatically synthesizes and validates unique inputs from given specification.
To cover the entire input space (and hence increase our chances to also cover the behavior and output spaces), we use grammar-based input producers (Zeller et al., 2023). These encode input specifications as context-free grammars (CFG) (Chomsky, 1956)—a finite set of recursive string generation rules. For some input formats, such CFGs already exist and are publicly available. For example, the Grammar Zoo (Zaytsev, 2015) repository contains definitions of 1756 grammars of mainstream programming languages mined from standard documentation or community-created resources. Developers could handwrite such grammars using the knowledge provided as protocol specifications such as RFCs. We cover alternative ways of automatically extracting grammars from existing systems in Section 6.10.
The Input Generator Module is the first component of the Modelizer framework. It acts as a data processing pipeline, as shown in Figure 3, consisting of four steps: (1) Synthesis, (2) Post-Processing, (3) Hashing, and (4) Validation. At the synthesis step, the input producer synthesizes new input samples according to rules encoded in grammar. Then, the synthesized samples are refined to better align with real-world data. Such refinement helps map synthesized input features to real-world data features without increasing grammar size, thereby simplifying the model learning process. The refined samples are then checked for uniqueness by computing a hash of the synthesized sample. If the hash is absent in the Hashes Dataset, the sample is passed to the PUT for further processing. Otherwise, all steps are repeated until a unique sample is generated. The PUT produces the corresponding output, which is saved in persistent storage with the input. The collected input-output pairs are later used to train a program behavior model.
@@
[…]
15<code_block> ::= ‘<text>‘ | ‘‘‘<text>‘‘‘ | <empty_space><text>
16<quoted_text> ::= ’<text>’ | "<text>"
17<empty_space> ::= " "
18<text> ::= TEXT
19<url> ::= URL
@@
Let us reuse our handwritten Markdown grammar (LABEL:lst:markdown-grammar) to demonstrate the synthetic input generation process. We will follow the steps that Modelizer takes to produce inputs. LABEL:lst:generate shows the high-level implementation of the input generation pipeline written in Python. Once the input producer is configured, it is ready to synthesize as many inputs as needed (Step 1). Here is an example of synthesized input that was generated by the producer:
@@
1def generate_data(grammar, min_non_term, max_non_term, total_elements, program_handler):
@@
5 fuzzer = GrammarFuzzer(grammar, min_non_term, max_non_term)
@@
The producer has synthesized one top-level Markdown element, a paragraph block. The paragraph block was expanded to a single paragraph text element, which contains a sequence of formatted text, hyperlink, and code block elements. One of the formatted text elements was further expanded to a bold text element. The synthesized input contains TEXT and URL placeholders instead of real character sequences. Instead of encoding the whole language vocabulary into our grammar and synthesizing natural language sentences, we have decided to define placeholder tokens to replace the real-world data. (We will elaborate more about placeholders in the follow-up section.)
@@ 2.1.2. Placeholders
Synthesizing data that is equivalent to real-world counterparts is a challenging task. As different programs can process the same input in various ways, the input grammar must be specified or refined according to program expectations. Depending on the complexity of the input specification, the number of required input samples for efficient behavior learning can also become extremely large. In their work on abstracting failure-inducing inputs, Gopinath et al. (Gopinath et al., 2020a) found that some input fragments are not affected or used by underlying implementation logic and can be abstracted for certain data types and programs. For example, many text format converters only process the document structure, leaving the document content untouched. To benefit from similar effects for behavior extraction, we integrated the following two strategies in our input generation pipeline:
@@
Both strategies will help capture large portions of possible input space with few samples and correctly map real-world inputs to generic representations and vice versa during the inference phase. Instead of learning the behavior from real-world examples, we learn how the PUT processes placeholders. (In the current Modelizer implementation, developers have to specify how to abstract grammar rules; future revisions will automate this process.)
@@
We assume that a TEXT placeholder can represent any character sequence in a real Markdown document. Similarly, a URL placeholder can substitute an arbitrary definition. Learning from the small number of synthesized structures, our models will still be capable of processing principally equivalent real-world inputs. While the input producer’s output is well-formed and passes parsing guards, aligning real-world Markdown documents with the synthesized structure is challenging because it is hard to distinguish the generated placeholders. We have implemented a post-processing phase in the input generation pipeline as a generic Placeholder Processor module. It initializes with a list of encoded placeholders like [TEXT, URL] and augments each placeholder with a unique identifier. Such identifiers enhance the accuracy of bidirectional translations, especially when elements need reordering.
Since grammar-based producers do not maintain a list of generated samples, we implemented a sample validation procedure. This involves computing an HMAC-SHA-384 (Hansen and 3rd, 2011) value (Step 3.1) for the synthesized sample and checking if it’s stored in the Hashes Dataset (Step 3.2). If absent, the dataset gets updated with a new record (Step 4), and the synthesized input is passed to the PUT for output computation. Otherwise, the sample is dropped, and input generation repeats until a unique sample is created. Input-output pairs are saved in persistent storage for later processing.
@@ 2.2. Dataset Pre-processing
Neural machine translation models are artificial intelligence models that use deep neural networks to learn dependencies between source and target language sentences. Their success in automated machine translation tasks inspired us to model program behavior as a sequence-to-sequence translation task. In principle, current neural machine translation architectures can process arbitrarily complex inputs, but the problem’s computational complexity scales quadratically with the input length. So, practical implementations partition data into token sequences and process them in fragments. Then, models learn how to generate sequences of tokens in a target language by mapping the cross-token dependencies from a source language to a target language. Similarly, the program input and output must be converted into sequences of tokens for processing. While text partitioning is well-studied in natural language processing  (Manning et al., 2008; Honnibal and Montani, 2017), tokenizing the program input and output could be a challenging task.